40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 4 00:32:24 2019, gautam, Update, ALS, More locking updates ALS_singleArm.pngALS_OOL_20191003.pdf
    Reply  Sat Oct 5 21:26:34 2019, gautam, Update, ALS, Y-end green alignment tweaked ALSY_alignment.pngALSY_OOL.pdf
       Reply  Sun Oct 6 15:23:27 2019, gautam, Update, ALS, Arm control using error signals achieved ALSY_20191006.pdftransitionIRALS.pngarms_ALS.pdf
Message ID: 14943     Entry time: Sat Oct 5 21:26:34 2019     In reply to: 14938     Reply to this: 14944
Author: gautam 
Type: Update 
Category: ALS 
Subject: Y-end green alignment tweaked 

Summary:

I improved the alignment of the green beam into the Y arm cavity.

  • GTRY went from ~0.2 to ~0.25, see Attachment #1.
  • This resulted in improvement of the Y arm ALS noise above 💯Hz by a factor of ~5, see Attachment #2.
  • I tried controlling the two arm cavities in the CARM/DARM basis using ALS error signals - but didn't manage to successfully execute this transition today - this will be the commissioning goal for the upcoming week.

Details:

  • I had to do the alignment by tweaking the steering mirrors at EY - the PZTs didn't give me anywhere near enough range.
  • While I was at EY, I tried moving the two MM lenses mounted on translation stages to try and improve the mode-matching into the arm cavity - wasn't successful, still see a bunch of bullseye modes when I toggle the shutter.
  • They EY green layout would benefit from a do-over (basically just copy the EX layout), but this isn't the priority right now, the ALS noise RMS is dominated by low frequency noise (as usual). 
  • There is a ~5% leakage of the GTRX beam onto the GTRY photodiode.
  • One thing to try would be to revive the MCL loop to reduce the <1 Hz laser frequency noise and see if that helps - basically testing this hypothesis.
  • I had done some careful noise-budgeting of the EX green PDH system, the EY system would benefit from the same, but not critical.
  • The improvement of the high-frequency noise is clear, and now we are consistent with the "known good reference" level from the time the DRFPMI locking was working back in early 2016.

Other changes made today:

  1. /opt/rtcds/caltech/c1/scripts/general/videoscripts/videoswitch was modified to be python3 compatible - for some reason, there were many syntax errors being thrown (even though I was using python2.7) and I wasn't able to change the displays in the VEA using the MEDM screen, but now it works again 👍.
  2. The LSC overview and several daughter MEDM screens were edited to remove references to channels that no longer exist. All screens I edited have a backup stored in the MEDM directory with today's date as a suffix.
  3. Input pointing into the PMC was tweaked.
  4. Noted that some pump is noisy at pumpspool - also noted that the annuli are no longer pumped. Some event seems to have triggered an interlock condition that closed off the annular volume from TP3, needs investigation...
Attachment 1: ALSY_alignment.png  166 kB  Uploaded Sat Oct 5 22:35:19 2019  | Hide | Hide all
ALSY_alignment.png
Attachment 2: ALSY_OOL.pdf  51 kB  Uploaded Sat Oct 5 22:35:25 2019  | Hide | Hide all
ALSY_OOL.pdf
ELOG V3.1.3-